OpenStudy (anonymous):

A stone is dropped from a tower 100 meters above the ground. The stone falls past ground level and into a well. It hits the water at the bottom of the well 5.00 seconds after being dropped from the tower. Calculate the depth of the well. Given: g = -9.81 meters/second2. 22.5 meters 50.7 meters 100 meters 110 meters 152.45 meters

OpenStudy (surry99):

What have you tried so far?

OpenStudy (anonymous):

22.5

OpenStudy (anonymous):

let the depth of the well is d Now , Using second equation of motion d = u * t + 0.5 g * t^2 100 + d = 0 * 5 + 0.5* 9.8 * 5^2 solving for d d = 22.5 m the depth of the well is 22.5 m